# Code-Embed: A Family of Open Large Language Models for Code Embedding

### v2

@@ Abstract
[…]
CodeXEmbed: A Generalist Embedding Model Family
for Multiligual and Multi-task Code Retrieval
Ye Liu, Rui Meng, Shafiq Joty, Silvio Savarese, Caiming Xiong, Yingbo Zhou, Semih Yavuz Salesforce AI Research yeliu@salesforce.com
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across numerous Natural Language Processing (NLP) tasks. However, they often struggle to produce faithful answers and may lack up-to-date or domain-specific knowledge. To bridge this gap, retrieval-augmented generation (RAG) Cai et al. (2022); Cheng et al. (2024) techniques have gained prominence, integrating Information Retrieval (IR) systems with LLMs to enhance their access to relevant external information. This synergy has drawn significant attention recently, leading to the development of various retrieval models Wang et al. (2022); Chen et al. (2024) based on BERT Kenton and Toutanova (2019) and other LLMs with sizes exceeding 1 billion parameters Wang et al. (2023); Moreira et al. (2024); Meng et al. (2024). Despite these advancements, standard IR methods, while effective in text-based retrieval, often fall short in specialized domains such as code retrieval Husain et al. (2019).
@@
Code retrieval is critical for accelerating development processes and improving code quality. Unlike general text retrieval, code retrieval enables developers to quickly locate relevant code snippets, explanations, bug analyses, summaries, and similar instances. Effective code retrieval systems are now integrated into commercial products like VS Code Del Sole (2021) and GitHub Copilot Wermelinger (2023); Yetiştiren et al. (2023), enhancing productivity. Code-RAG systems Parvez et al. (2021); Liu et al. ; Jimenez et al. (2024); Wang et al. (2024) also leverage code retrieval to minimize hallucinations in generated code from LLMs, ensuring more accurate outputs. However, traditional text retrieval models often struggle with code because they focus on linguistic patterns, while code retrieval must handle elements like syntax, variable dependencies, control flow, and API usage. Despite the importance of code-specific models, existing ones like CodeBERT Feng et al. (2020), CodeGPT Lu et al. , and UniXcoder Guo et al. (2022) are based on smaller BERT models Kenton and Toutanova (2019). While large-scale models for retrieval have become popular, only Voyage-Code has followed this approach for code retrieval, but it remains a closed model, leaving a gap for open-source, large-scale code retrieval models.
@@
In this work, we introduce CodeXEmbed, a family of open-source embedding models tailored for both code and text, available in sizes of 400 million, 2 billion, and 7 billion parameters. CodeXEmbed introduces a generalizable training framework that converts multiple programming languages and diverse code-related tasks into retrieval tasks. Our approach handles 12 programming languages and five distinct code retrieval categories across eight different code tasks, including code-to-text, text-to-code, code-to-code, text-to-text and hybrid text and code tasks. This comprehensive setup enables CodeXEmbed to generalize effectively across various code domains. Our contributions can be summarized as follows:
@@
  • We propose a generalizable training approach for code embedding that transforms diverse programming languages and code-related tasks into retrieval tasks. This method allows for a unified framework that effectively handles multiple languages and task types, improving retrieval performance across various code tasks.
@@
  • Our 7B model, the open-source code embedding model to date, achieves state-of-the-art performance on code retrieval CoIR benchmark, setting a new standard for code retrieval. We also evaluate CodeXEmbed on RepoEval and SWE-Bench-Lite, demonstrating that improved retrieval significantly enhances end-to-end Retrieval-Augmented Generation (RAG) performance for code tasks.
  • In addition to the 7B model, we train smaller models (400M and 2B) that outperform the previous state-of-the-art in code retrieval. Furthermore, our models demonstrate competitive performance on text retrieval, making them versatile across both code and text retrieval tasks.
@@ 2 Method
We transform general code-related tasks into a unified retrieval framework by categorizing them into distinct retrieval settings. As shown in Figure 1, each setting serves a specific purpose in mapping between code and text, enhancing the model’s ability to generalize across various retrieval tasks.
@@
Text-to-Code Retrieval. The Text-to-Code retrieval setting focuses on retrieving relevant code documents based on a given textual query. This setup is particularly useful in scenarios where natural language descriptions are mapped to code solutions. Several code generation tasks can be transformed to fit into this retrieval paradigm.
One such task is code contest generation Billah et al. (2024); Kadir et al. (2024), where a natural language description of a programming problem is provided, and the goal is to retrieve or generate a code snippet that correctly solves the problem. This setting closely aligns with the Text-to-Code retrieval framework as it involves finding code solutions based on a question or problem statement. Another example is the Text-to-SQL task Finegan-Dollak et al. (2018); Li et al. (2024a), where the aim is to generate or retrieve a SQL query based on a user’s natural language question. The task fits well within the Text-to-Code category, as it requires understanding the intent of the query and retrieving or generating the appropriate SQL code that matches the given description.
Code-to-Text Retrieval. The Code-to-Text retrieval setting is designed to retrieve relevant textual documents or descriptions based on a given code query. This setup is particularly valuable in scenarios where developers or systems seek to understand or explain code by mapping it to human-readable documentation or summaries.
One such task is code summarization Sontakke et al. (2022); Sun et al. (2024), where a code file, or even an entire repository, is provided as input, and the goal is to generate a concise textual summary of the code’s functionality. This task aligns with the Code-to-Text retrieval setting, as it involves taking potentially complex code and producing a human-readable description or explanation, which is critical for documentation, knowledge sharing, and onboarding in software projects.
Code-to-Code Retrieval. The Code-to-Code retrieval setting focuses on retrieving relevant code documents based on a code query. This retrieval setting is particularly useful for tasks that involve understanding, transforming, or complementing code snippets. Several tasks naturally align with this framework, making it an essential part of the broader code retrieval ecosystem.
One such task is code translation Pan et al. (2024); Karanjai et al. (2024), where the goal is to translate code snippets from one programming language to another. For instance, a developer might provide a code snippet in Python and seek to translate it into C++ or Java. This task fits well into the Code-to-Code retrieval setting, as it involves finding the equivalent functionality in a different programming language, ensuring the translated code maintains the same logic and structure.
Another important task is code completion Ding et al. (2024); Phan et al. (2024); Liu et al. (2024), where a partially written code snippet is provided, and the system needs to retrieve or generate the missing portions to complete the function or module. This is commonly seen in integrated development environments (IDEs) where autocomplete features assist developers by suggesting the next logical lines of code based on the current context. Code completion aligns perfectly with the Code-to-Code retrieval setting, as it requires matching the partial input code with relevant snippets that seamlessly complement the existing code.
Additionally, code clone detection Martinez-Gil (2024), which involves identifying code snippets that are functionally or structurally identical but may differ in syntax or formatting, also fits within this retrieval framework. Clone detection is crucial for identifying redundant code, refactoring codebases, and ensuring consistency across different sections of a project. This task leverages the Code-to-Code retrieval setting by retrieving code snippets that exhibit similar or identical behavior, even if they appear different at the surface level.
Hybrid Code Retrieval. The hybrid code retrieval setting involves retrieving a combination of both code and textual documents in response to a hybrid query, which may contain both code and natural language. This setting is particularly useful for tasks that require understanding and integrating both code and textual context to generate meaningful responses or retrieve relevant documents.
One such task is code agent conversation Arteaga Garcia et al. (2024); Jin et al. (2024), where a code agent must generate either a code snippet or a textual response based on a user’s question. This task requires the agent to interpret the user’s question, which might contain both code and natural language, and retrieve or generate the appropriate code solution or explanation. The hybrid query allows the agent to navigate between technical (code) and descriptive (textual) domains to provide a comprehensive response.
Another key task is code issue fixing Yang et al. (2024); Jimenez et al. (2024), where a hybrid query containing a description of the code issue and the problematic code snippet is provided. The goal is to generate code that resolves the issue. By incorporating both the textual description (e.g., error message) and the code itself, hybrid retrieval allows the model to identify relevant code snippets, detect similar issues, and propose fixes that align with the overall project. This makes hybrid retrieval essential in complex development environments, where both debugging and problem-solving require a nuanced understanding of both code and text.
Text-to-Text Retrieval. In order to capture the model’s general retrieval ability across both code and natural language domains, we incorporate text-to-text retrieval data into the training pipeline. This allows the model to perform well in traditional text retrieval tasks, enhancing its versatility. The experiments study the ratio of text and code data to determine the optimal balance between the two, aiming to ensure the model’s effectiveness in both text and code retrieval tasks without sacrificing performance in either domain.
Retrieval Training. To train a unified code and text retrieval model, we transform all the aforementioned tasks into a query-and-answer format, where either the query or the answer can be a text or code document. This unified format allows us to handle a variety of tasks such as Text-to-Code, Code-to-Text, and Code-to-Code retrieval within the same training framework. The training loss is designed as a contrastive loss Liu et al. (2021), which aims to maximize the similarity between the query and the correct answer while minimizing the similarity to the negative samples. The loss function is represented as:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}\log\frac{\exp(q_{i}\cdot d_{i}^{+})}{\exp(q_{i}\cdot d_{i}^{+})+\sum_{j=1}^{K}\exp(q_{i}\cdot d_{j}^{-})}\Bigg{]}$ (1)
where $q_{i}$ represents the query, which can be either text or code. The term $d_{i}^{+}$ refers to the corresponding positive document, which could also be in the form of text or code. On the other hand, $d_{j}^{-}$ denotes a set of hard negative documents, which are selected to be similar in context to the query but incorrect, making the retrieval task more challenging. The similarity between the query and a document is measured by $\text{sim}(q,d)$. The variable $N$ represents the total number of training samples, indicating the number of queries used during training, while $K$ defines the number of hard negatives associated with each query.
@@ 3 Experiments
Evaluation Benchmarks. We mainly use two benchmarks to evaluate code and text retrieval performance. COIR Li et al. (2024b) is a comprehensive benchmark specifically designed for code retrieval tasks. COIR covers a wide range of retrieval challenges, including 8 fine-grained retrieval subtasks, spanning 14 major programming languages. The dataset is composed of 10 distinct datasets, with a total corpus exceeding 2 million entries. BEIR Thakur et al. is a widely-adopted benchmark designed for text retrieval tasks. BEIR encompasses a diverse set of retrieval challenges, covering 9 distinct tasks across various domains such as question answering, duplicate detection, fact-checking, and more. It supports retrieval over a wide range of datasets and provides a standardized benchmark for evaluating text retrieval models across different domains.
@@
Implementation Details. We conduct general training on our proposed code and text pair dataset using three model sizes: 400M, 2B, and 7B. For the CodeXEmbed400M, we use the base model Alibaba-NLP/gte-large-en-v1.5 (Li et al., 2023b), applying full model fine-tuning. For the CodeXEmbed2B, we initialize our embedding model from the generation model google/gemma-2-2b-it Team et al. (2024), using low-rank adaption(LoRA) Hu et al. (2022) with a rank of 8. For the CodeXEmbed7B, we initialize our embedding model from the generation model mistralai/Mistral-7B-Instruct-v0.3, also using LoRA with a rank of 8. Following prior work Meng et al. (2024), we apply: (i) last token pooling for CodeXEmbed2B and CodeXEmbed7B, and (ii) beginning token pooling for CodeXEmbed400M to generate semantic vector representations Li et al. (2023b). Cosine similarity is used to compute the similarity between query and corpus for ranking. The batch size is set to 1024 across all three model sizes, with 7 hard negatives. The learning rate is set to $5e^{-5}$, and the end learning rate to $5e^{-6}$, with linear decay and a 50-step warmup. To improve training efficiency and reduce GPU memory usage, we adopt gradient caching Gao et al. (2021). The more implementation details can be found in Appendix A.2.
@@
Baseline Models. For code-domain-specific models, we included UniXcoder (Guo et al., 2022) and Voyage-Code-002111https://blog.voyageai.com/2024/01/23/voyage-code-2-elevate-your-code-retrieval/, both of which are pre-trained on code data, making them strong baselines for comparison. For general retrieval models, we evaluated E5-Base (Wang et al., 2022), GTE-Base (Li et al., 2023b), BGE-Base (Xiao et al., 2023), Contriever (Izacard et al., 2023), E5-Mistral (Wang et al., 2023), BGE-M3 (Chen et al., 2024), NV-Embed-V2 Moreira et al. (2024), SFR-V2 Meng* et al. (2024) and OpenAI-Ada-002222https://platform.openai.com/docs/guides/embeddings.
@@
Evaluation Metrics. In code retrieval, selecting the right evaluation metric is key for assessing both ranking sensitivity and relevance. Building on prior work Wang et al. (2013), Normalized Discounted Cumulative Gain (NDCG) is preferred for its ability to account for both rank order and varying relevance, unlike precision, recall or MRR. Therefore, we use NDCG@10 to evaluate performance on CoIR and BEIR.

### v3

# CodeXEmbed: A Generalist Embedding Model Family for Multilingual and Multi-task Code Retrieval

@@ Abstract
Despite the success of text retrieval in many NLP tasks, code retrieval has received comparatively less attention than text retrieval. Most retrieval systems are designed for natural language queries and often fail to address the structural and semantic complexities of retrieving code. Consequently, they struggle across diverse programming languages and retrieval tasks, underscoring the need for specialized models tailored to code retrieval. To address this gap, we introduce CodeXEmbed, a family of large-scale code embedding models ranging from 400M to 7B parameters. Our novel training pipeline integrates multiple programming languages and reformulates diverse code-related tasks within a unified retrieval framework, enhancing model generalizability and performance. Our largest model (7B parameters) sets a new state-of-the-art (SOTA) in code retrieval, ranking first on the CoIR Leaderboard. Beyond code retrieval, our models achieve competitive results on the widely adopted BeIR text retrieval benchmark, demonstrating cross-domain versatility. Furthermore, our findings highlight that advancements in retrieval quality directly improve end-to-end Retrieval-Augmented Generation (RAG) performance for code-related tasks.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across numerous Natural Language Processing (NLP) tasks. However, they often struggle to produce faithful answers and may lack up-to-date or domain-specific knowledge. To bridge this gap, retrieval-augmented generation (RAG) Cai et al. (2022); Cheng et al. (2024); Nguyen et al. (2024) techniques have gained prominence, integrating Information Retrieval (IR) systems with LLMs to enhance their access to relevant external information. This synergy has drawn significant attention recently, leading to the development of various retrieval models Wang et al. (2022); Chen et al. (2024) based on BERT Kenton & Toutanova (2019) and other LLMs with sizes exceeding 1 billion parameters Wang et al. (2023); Moreira et al. (2024); Meng et al. (2024). Despite these advancements, standard IR methods, while effective in text-based retrieval, often fall short in specialized domains such as code retrieval Husain et al. (2019); Li et al. (2024b).
@@
Code retrieval accelerates software development and improves code quality by enabling quick access to relevant snippets, explanations, and analyses. Unlike general text retrieval, it must handle syntax, dependencies, and control flow. Integrated into tools like VS Code Del Sole (2021) and GitHub Copilot Wermelinger (2023); Yetiştiren et al. (2023), code retrieval also enhances Code-RAG systems Parvez et al. (2021); Liu et al. (2020); Jimenez et al. (2024); Wang et al. (2024) by reducing LLM hallucinations. However, traditional text retrieval models struggle with code-specific elements. Existing models like CodeBERT Feng et al. (2020), CodeGPT Lu et al. (2021), and UniXcoder Guo et al. (2022) are based on smaller BERT models Kenton & Toutanova (2019) leading to subpar performance. Furthermore, a unified model capable of handling both text and code retrieval is essential for seamless integration, enabling developers to retrieve documentation, explanations, and relevant code within a single framework.
@@
In this work, we introduce CodeXEmbed  [^1] Model weights: https://huggingface.co/Salesforce/SFR-Embedding-Code-2B_R, https://huggingface.co/Salesforce/SFR-Embedding-Code-400M_R., a family of open-source embedding models tailored for both code and text, available in sizes of 400 million, 2 billion, and 7 billion parameters. CodeXEmbed introduces a generalizable approach that converts diverse code-related tasks with various programming languages into a unified contrastive training framework. Our approach handles 12 programming languages and five distinct code retrieval categories across eight different code tasks, including code-to-text, text-to-code, code-to-code, text-to-text and hybrid text and code tasks. To enable the model to retrieve both text and code effectively, we propose a novel multi-stage training method. In the first stage, we train the model on text data using LoRA adaptation. In the second stage, we continue training with both code and text data while tuning only the newly introduced LoRA adapter. This approach preserves the text retrieval knowledge learned in the first stage while enhancing code retrieval, ensuring the model retains strong performance across both domains. This comprehensive setup enables CodeXEmbed to generalize effectively across various code domains. Our contributions can be summarized as follows:
@@
  • We introduce a generalizable multi-stage training approach for code and text embedding, unifying diverse code-related tasks within a retrieval framework. This leads to significant improvements in retrieval performance across multiple programming languages and tasks.
@@
  • Our 7B model achieves state-of-the-art performance on the CoIR benchmark, setting a new standard for code retrieval. We furhter evaluate CodeXEmbed on RepoEval and SWE-Bench-Lite, demonstrating that improved retrieval significantly enhances end-to-end Retrieval-Augmented Generation (RAG) performance for code-related tasks.
  • Beyond the 7B model, we develop smaller models (400M, 2B) that surpass prior SOTA in code retrieval, while remaining competitive performance in text retrieval, highlighting their versatility across both domains.
@@ 2 Method
We transform general code-related tasks into a unified retrieval framework by representing each task as a structured retrieval pair $(Q,D)$, where $Q$ is the query and $D$ is the positive document. The retrieval process is formulated as a ranking problem, where the objective is to learn a function $f(Q,D)$ that assigns higher similarity scores to relevant pairs than that of irrelevant ones.
2.1 Unified Retrieval Framework
To train a unified retrieval model for both code and text, we transform all retrieval tasks into a query-document matching problem, where a query $Q$ retrieves a relevant document $D$ from a candidate set $\mathcal{D}$. This formulation supports Text-to-Text, Text-to-Code, Code-to-Text, and Code-to-Code retrieval within a single framework.
To better capture semantic relationships across entire input sequences, we incorporate bidirectional attention into a causal LLM BehnamGhader et al. (2024), enhancing its ability to encode contextual information for retrieval tasks. Each query $Q$ and document $D$ are encoded using this pre-trained LLM, producing representations:
$q=\text{LLM}_{\theta+\Delta\theta}(Q),\quad d=\text{LLM}_{\theta+\Delta\theta}(D)$
The similarity between query and document is measured using cosine similarity:
$D^{*}=\arg\max_{D\in\mathcal{D}}\text{sim}(q,d),\quad\text{sim}(q,d)=\frac{q\cdot d}{\|q\|\|d\|}$ (1)
2.2 Multi-Stage Training with LoRA
To progressively enhance retrieval capabilities, we adopt a multi-stage training approach, utilizing distinct LoRA Hu et al. (2022) adapters at each stage. In each stage, the previously trained LoRA weights are merged into the base model before initializing a new adapter, as shown in Figure 1. This approach enables the model to iteratively adapt to different retrieval objectives without catastrophic forgetting Magistri et al. (2024); Han et al. (2024)
Figure 1: The illustration depicts multi-stage training, showing the data used at each stage, the LoRA adapters applied, and the continuous optimizer with its learning rate progression.
2.2.1 Stage 1: Supervised Contrastive Pretraining on Text Data
We initialize a LoRA adapter $\Delta\theta_{1}$ and freeze the LLM $\theta$ and train $\Delta\theta_{1}$ on text-based retrieval tasks using a contrastive learning objective:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(q_{i},d_{i}^{+}))}{\exp(\text{sim}(q_{i},d_{i}^{+}))+\sum_{j=1}^{K}\exp(\text{sim}(q_{i},d_{j}^{-}))}$ (2)
where $d_{i}^{+}$ is a positive document, and $d_{j}^{-}$ represents negative documents, and $N$ denotes the batch size. The model learns to align text queries with their corresponding text-based documents, establishing a strong foundation for language-based retrieval.
2.2.2 Stage 2: Code-Enhanced Retrieval Training
After training, we merge the trained adapter $\Delta\theta_{1}$ into the base LLM model:
$\theta\leftarrow\theta+\Delta\theta_{1}$ (3)
and initialize a new LoRA adapter $\Delta\theta_{2}$ to learn retrieval for both code and text. This setup enables the model to mitigate text-only biases while preserving most of its learned representations.
To ensure a smooth transition between training phases, we continue optimizing the second-stage LoRA adapter $\Delta\theta_{2}$ while retaining the gradient contributions from the first-stage adapter $\Delta\theta_{1}$, which remains fixed within the model. Specifically, we update $\Delta\theta_{2}$ while preserving the learning rate $r^{(1)}$ and the accumulated gradient from $\Delta\theta_{1}$:
$\Delta\theta_{2}\leftarrow\text{Update}(\Delta\theta_{2},\nabla L(\theta_{1},\theta_{2}),r^{(1)})$ (4)
where $\nabla L(\theta_{1},\theta_{2})$ represents the gradient influenced by both adapters, but only $\Delta\theta_{2}$ is actively updated. This approach stabilizes optimization, preventing abrupt parameter shifts while allowing a smooth adaptation to code retrieval. By preserving the previous learning rate and gradient continuity, we mitigate convergence instability and facilitate efficient adaptation.
This multi-stage approach enables the model to iteratively improve retrieval performance while preventing overfitting to any single modality. By leveraging LoRA adapters at different stages and maintaining a smooth learning rate transition, we ensure efficient adaptation with minimal overhead, resulting in more robust text and code retrieval.
2.3 Scaling Batch Size with GradCache
Larger batch sizes improve contrastive learning by incorporating more diverse negatives Chen et al. (2022), but GPU memory limits expansion. To overcome this, we use GradCache Gao et al. (2021), which reduces memory overhead, enabling larger effective batch sizes. Let $Q/D$ denote the full batch of queries and documents, partitioned into sub-batches ($Q=\{\hat{Q}_{1},\hat{Q}_{2},\dots,\hat{Q}_{M}\}$, $D=\{\hat{D}_{1},\hat{D}_{2},\dots,\hat{D}_{M}\}$) to fit memory constraints. Training follows three steps:
Graph-less Forward Pass Each sub-batch is encoded into embeddings $q_{i}$ and $d_{i}$ without computing encoder gradients, reducing memory usage.
Gradient Computation and Caching For each query $q_{i}$ in sub-batch $\hat{Q}_{j}$, we compute and cache gradients w.r.t. the representation function:
$\mathbf{u}_{i}=\frac{\partial\mathcal{L}}{\partial q_{i}},\quad\mathbf{v}_{i}=\frac{\partial\mathcal{L}}{\partial d_{i}}$ (5)
Sub-batch Gradient Accumulation We reprocess sub-batches to collect embeddings $q_{i}$ and $d_{i}$, construct the computation graph, and link gradients:
$\frac{\partial\mathcal{L}}{\partial\theta}=\sum_{\hat{Q}_{j}\in Q}\sum_{q_{i}\in\hat{Q}_{j}}\mathbf{u}_{i}\frac{\partial q_{i}}{\partial\theta}+\sum_{\hat{D}_{j}\in D}\sum_{d_{i}\in\hat{D}_{j}}\mathbf{v}_{i}\frac{\partial d_{i}}{\partial\theta}$ (6)
$\partial\theta$ denotes the LoRA adapter in either stage 1 or 2. GradCache enables larger in-batch negatives, improving contrastive learning efficiency while staying within GPU limits, essential for scaling text and code retrieval.
@@ 2.3 Scaling Batch Size with GradCache
2.4 Unified Retrieval Training Data
We organize retrieval tasks into distinct settings by converting various code-related tasks into a unified retrieval format. In this format, the input is treated as query ($Q$), and the expected output or relevant content is treated as document ($D$). This transformation allows the model to generalize across various retrieval tasks in both text and code domains.
2.4.1 Text Retrieval Training Data
To improve retrieval across both code and text domains, we incorporate text-to-text retrieval tasks from the BEIR datasets Thakur et al. (2021). This ensures model robustness in general document retrieval while maintaining strong performance in code retrieval.
2.4.2 Code Retrieval Training Data
We unify multiple code generation and classification tasks into the $(Q,D)$ format for code retrieval training, as shown in Figure 2. Text-to-Code Retrieval involves retrieving relevant code snippets given a natural language query. Tasks such as code contest generation Billah et al. (2024); Kadir et al. (2024) take problem descriptions as $Q$ and map them to the correct implementation as $D$. Similarly, in Text-to-SQL Finegan-Dollak et al. (2018); Li et al. (2024a), a user query in natural language serves as $Q$, and the corresponding SQL query is treated as $D$. Code-to-Text Retrieval retrieves human-readable descriptions for given code snippets. In code summarization Sontakke et al. (2022); Sun et al. (2024), a code snippet serves as $Q$, and the corresponding documentation or summary is $D$. Code-to-Code Retrieval represents retrieving functionally equivalent or semantically related code snippets. In code translation Pan et al. (2024); Karanjai et al. (2024), the source code in one language is $Q$, and its equivalent implementation in another language is $D$. Code completion Ding et al. (2024); Phan et al. (2024); Liu et al. (2024) uses the incomplete code fragment as $Q$ and the correct completion as $D$. Code clone detection Martinez-Gil (2024) retrieves functionally similar code, treating the reference snippet as $Q$ and the detected duplicate as $D$. Hybrid Code Retrieval supports mixed queries containing both text and code. In code agent conversation Arteaga Garcia et al. (2024); Jin et al. (2024), user prompts containing explanations or code snippets act as $Q$, while the retrieved response (either code or text) serves as $D$. In code issue fixing Yang et al. (2024); Jimenez et al. (2024), the hybrid query consists of an error message and buggy code as $Q$, while the corrected version of the code is $D$.
By consolidating diverse retrieval tasks into a unified framework, we effectively streamline contrastive model training, thereby facilitating seamless adaptation across a wide range of programming scenarios while ensuring high retrieval accuracy.
@@ 3 Experiments
Evaluation Benchmarks We mainly use two benchmarks to evaluate code and text retrieval performance. COIR Li et al. (2024b) is a comprehensive benchmark specifically designed for code retrieval tasks. COIR covers a wide range of retrieval challenges, including 8 fine-grained retrieval subtasks, spanning 14 major programming languages. The dataset is composed of 10 distinct datasets, with a total corpus exceeding 2 million entries. BEIR Thakur et al. (2021) is a widely-adopted benchmark designed for text retrieval tasks. BEIR encompasses a diverse set of retrieval challenges, covering 9 distinct tasks across various domains such as question answering, duplicate detection, fact-checking, and more. It supports retrieval over a wide range of datasets and provides a standardized benchmark for evaluating text retrieval models across different domains.
@@
Implementation Details We conduct general training on our proposed code and text pair dataset using three model sizes: 400M, 2B, and 7B. For the CodeXEmbed400M, we use the base model Alibaba-NLP/gte-large-en-v1.5 (Li et al., 2023b), applying full model fine-tuning. For the CodeXEmbed2B, we initialize our embedding model from the generation model google/gemma-2-2b-it Team et al. (2024), using low-rank adaption(LoRA) Hu et al. (2022) with a rank of 8. For the CodeXEmbed7B, we initialize our embedding model from the generation model mistralai/Mistral-7B-Instruct-v0.3, also using LoRA with a rank of 8. Following prior work Meng et al. (2024), we apply: (i) last token pooling for CodeXEmbed2B and CodeXEmbed7B, and (ii) beginning token pooling for CodeXEmbed400M to generate semantic vector representations Li et al. (2023b). Cosine similarity is used to compute the similarity between query and corpus for ranking. The batch size is set to 1024 across all three model sizes, with 7 hard negatives. The learning rate is set to $5e^{-5}$, and the end learning rate to $5e^{-6}$, with linear decay and a 50-step warmup. To improve training efficiency and reduce GPU memory usage, we adopt gradient caching Gao et al. (2021). The more implementation details can be found in Appendix A.3.
@@
Baseline Models For code-domain-specific models, we included UniXcoder (Guo et al., 2022), Voyage-Code-002222https://blog.voyageai.com/2024/01/23/voyage-code-2-elevate-your-code-retrieval/ and CodeSage-large-v2 (Zhang et al., 2024), all pre-trained on code data, serving as strong baselines for comparison. For general retrieval models, we evaluated E5-Base (Wang et al., 2022), GTE-Base (Li et al., 2023b), BGE-Base (Xiao et al., 2023), Contriever (Izacard et al., 2023), E5-Mistral (Wang et al., 2023), BGE-M3 (Chen et al., 2024), NV-Embed-V2 Moreira et al. (2024), SFR-V2 Meng* et al. (2024) and OpenAI-Ada-002333https://platform.openai.com/docs/guides/embeddings.
@@
Evaluation Metrics In code retrieval, selecting the right metric is key for assessing both ranking sensitivity and relevance. Following prior work Wang et al. (2013), Normalized Discounted Cumulative Gain (NDCG) is preferred for its ability to account for both rank order and varying relevance. Therefore, we use NDCG@10 to evaluate performance on CoIR [^4] CoIR Implementation https://github.com/CoIR-team/coir and BEIR.